Lorenta Company has provided the following data (ignore income taxes):

2019 revenues were $78,700.

2019 expenses were $50,300.

Dividends declared and paid during 2019 totaled $9,400.

Total assets at December 31, 2019 were $202,000.

Total liabilities at December 31, 2019 were $139,000.

Common stock at December 31, 2019 was $44,000.

Which of the following is not correct?

Multiple Choice

  • 2019 net income was $28,400. (A)

  • Total stockholders' equity at December 31, 2019 was $63,000. (B)

  • Total liabilities and stockholders' equity at December 31, 2019 was $202,000. (C)

  • Retained earnings on December 31, 2019 were $202,000 (D)
Answer #1

Net income = Revenues - Expenses = 78,700 - 50,300 = 28,400

Total stockholders equity = Total assets - Total Liabilities

= 202,000 - 139,000 = 63,000

Total Liabilities and stockholders equity = Total assets = 202000

Retained earnings = Net income - Dividends paid

= 28,400 - 9,400

= 19,000 (Option D is the answer)